# MMGPL: Multimodal Medical Data Analysis with Graph Prompt Learning

### v1

@@ Abstract
Prompt learning has demonstrated impressive efficacy in the fine-tuning of multimodal large models to a wide range of downstream tasks. Nonetheless, applying existing prompt learning methods for the diagnosis of neurological disorder still suffers from two issues: (i) existing methods typically treat all patches equally, despite the fact that only a small number of patches in neuroimaging are relevant to the disease, and (ii) they ignore the structural information inherent in the brain connection network which is crucial for understanding and diagnosing neurological disorders. To tackle these issues, we introduce a novel prompt learning model by learning graph prompts during the fine-tuning process of multimodal large models for diagnosing neurological disorders. Specifically, we first leverage GPT-4 to obtain relevant disease concepts and compute semantic similarity between these concepts and all patches. Secondly, we reduce the weight of irrelevant patches according to the semantic similarity between each patch and disease-related concepts. Moreover, we construct a graph among tokens based on these concepts and employ a graph convolutional network layer to extract the structural information of the graph, which is used to prompt the pre-trained multimodal large models for diagnosing neurological disorders. Extensive experiments demonstrate that our method achieves superior performance for neurological disorder diagnosis compared with state-of-the-art methods and validated by clinicians.
@@ I Introduction
Figure 1: The flowchart of the proposed MMGPL consisting of three modules i.e., multimodal data tokenizer (light blue block), concept learning (light green block), and graph prompt learning (light yellow block). First, we first divide the raw multimodal data into multiple patches and project them into a shared embedding space (Sec. III-B). Second, we prompt the GPT-4 to generate disease-related concepts and further learn the weights of tokens based on the semantic similarity between tokens and concepts (Sec. III-C). Third, We learn a graph among tokens and extract structural information to prompt the unified encoder (Sec. III-D). Finally, we obtain the output from the unified encoder and use it to predict the label of the subject.
Firstly, only a small fraction of patches in neural data are related to the disease. Unlike natural image data, the pathological regions or the regions of interest (i.e., ROIs) in neural data typically only occupy a small portion of the entire data. For example, previous studies (Padurariu et al., 2012; Wang et al., 2006) have shown a significant correlation between the hippocampal area and Alzheimer’s disease. Recently, in the area of medical image segmentation, methods (Huang et al., 2023; Zhang et al., 2023b) utilize region-based prompts, such as bounding boxes and clicks, to guide the fundamental model focus on ROIs within the medical images. Despite the promising results obtained by region-based prompts, the aforementioned operations (e.g., bounding box and clicks) often rely on human interaction or object detection models which reduces the flexibility of these methods. Moreover, implementing the aforementioned operations becomes even more challenging when dealing with multimodality neural data, which may typically contains 3D tensor MRI data, time-series data, and functional connectivity (FC) networks data. Hence, it is challenging to discover useful patches when dealing with multimodality neural data.
@@ II-A Multimodal large models
Unlike traditional large language models such as GPT-3 (Brown et al., 2020) and BERT (Devlin et al., 2018), which can only deal with text data, multimodal large models (Radford et al., 2021; Tsimpoukelli et al., 2021; Liu et al., 2023; Driess et al., 2023) can integrate multiple modalities, such as language, vision, and audio, to perform a broad spectrum of tasks. For instance, CLIP (Radford et al., 2021) leverages contrastive learning approach for joint pre-training on image-text pairs data. Besides, LLaVA (Liu et al., 2023) conduct image-text representation alignment followed by instruction tuning. Frozen (Tsimpoukelli et al., 2021) firstly proposed a approach that employing large language models in multimodal in-context learning. Similarly, Video-ChatGPT (Maaz et al., 2023) merges a video-adapted visual encoder with a large language models to handled video data. Recent works like Imagebind (Girdhar et al., 2023) and Meta-transformer (Zhang et al., 2023a) can learn a joint embedding across more than six different modalities.
In the medical domain, both large language models and multimodal large models have been recently explored for a wide range of medical tasks, including medical question answering and segmentation. For example, LLaVA-Med (Li et al., 2023) trains a vision-language conversational assistant research questions about medical images. MedSAM (Ma and Wang, 2023) fine-tune the segment anything model (SAM) (Kirillov et al., 2023) on medical image data for segmentation task. Moreover, Med-PaLM (Tu et al., 2023) and RedFM (Wu et al., 2023) interpret multimodal biomedical data and handle a diverse range of tasks, moving closer to generalist medical artificial intelligence models. However, due to the complexity of multimodal neuroimaging data, there is still a lack of research and remains challenges in leveraging multimodal large models for diagnosing neurological disorders.
@@ II-B Prompt learning
Prompt learning has emerged as a new paradigm in fine-turning pretrained large models for various downstream tasks. Compared to full fine-tuning, it achieves comparable performance without updating all parameters in large large models. In the field of NLP, GPT-3 (Brown et al., 2020) has demonstrated strong generalization to downstream tasks by manually selecting prompt texts. To improve the construction of prompt texts, Prefix-tuning (Li and Liang, 2021) optimizes prompts through gradient-based fine-tuning. In computer vision, VPT (Jia et al., 2022) prompts the image for effectively using the pre-trained image encoder in classification task. SAM (Kirillov et al., 2023) take visual prompts by drawing points, boxes, and strokes on an image for image segmentation. Moreover, MaPLe (Khattak et al., 2023) utilizes multi-modal prompts to guide models in both image and text modalities.
In the field of medical, prompt learning aims to guide models by focusing their attention on relevant regions within complex medical datasets. For example, several researchers (Ma and Wang, 2023; Huang et al., 2023; Zhang et al., 2023b) have introduced prompts such as bounding boxes and clicks into the SAM model to adapt it for medical image segmentation tasks. However, these methods may not be suitable for diagnosing neurological disorders given that they rely on human interaction or object detection models, as well as they do not take into account the importance of brain connectivity in neurological disorders.
@@ II-C Graph neural network
In the past decade, Graph Neural Networks (GNNs) have widely used in computer-aided diagnosis (Sun et al., 2020; Bessadok et al., 2022; Holzinger et al., 2021; Li et al., 2021). GNNs such as Graph Convolutional Network (GCN, Kipf and Welling (2017)) and Graph Attention Network (GAT) (Veličković et al., 2017) leverage message passing mechanism to capture relationships between nodes and structure information in the graph. Based on this, GNNs have the capability to leverage the inherent relationships between brain regions to uncover patterns or biomarkers. As a result, there is increasing interest in the application of GNNs in the field of neuroscience (Li et al., 2021). For example, Parisot et al. (2018) exploit the GCN for Alzheimer’s disease diagnose in population graph where the nodes in the graph denote the subjects and edges represent the similarity between subjects. Similarly, Kazi et al. (2019) propose a GCN model with multiple filter kernel sizes on a population graph. Moreover, In addition, brain graph is also frequently used in the diagnosis of neurological disorders, where the nodes correspond to anatomical brain regions and the edges depict functional or structural connections between these regions. For instance, Li et al. (2021) conduct an interpretable model on brain graph to determine the specific brain regions that are associated with a particular neurological disorder. Cai et al. (2022) propose a transformer-based geometric learning approach to handle multimodal brain graphs for brain age estimation. However, there is still a lack of research on how to effectively leverage graph in multimodal large models.
@@ III-A Preliminary and motivations
where LN is Layer normalization operation, $\mathbf{Z}^{(l-1)}$ is the output from the previous layer (and the input token embeddings for $l=1$). In the MHSA layer, the attention mechanism computes multiple attention scores by performing dot products between the Query and Key vectors of the tokens, which represent dense interactions among tokens. A single-head attention operation can be expressed as follows:
@@
$\mathbf{P}_{\text{combined}}=\mathcal{F}_{\text{conbine}}(\mathbf{P}_{\text{prompt}},\mathbf{P}),$ (4)
where $\mathcal{F}_{\text{conbine}}$ denotes a fusion layer that can be a non-parametric operator ($\oplus$ or $+$) or parameterized layers. The combined token embeddings $\mathbf{P}_{\text{combined}}$ are subsequently processed by the transformer encoder, as delineated in Eq. 2. Based on this, previous methods proposed cross-modality prompt learning (Khattak et al., 2023) and deep prompt learning (Liu et al., 2022) to further improve its performance on specific tasks
@@ III-B Multimodal data tokenizer
Medical data is inherently multimodal, usually including modalities like MRI, PET, and FC in neurological disorder diagnosis. These diverse modalities provide complementary information, helping to gain a more comprehensive understanding and better analyze neurological disorders. However, these modalities often exhibit more complex data structures compared to natural images, such as 3D tensor medical data, brain connectivity graphs data, and time series data. Previous multimodal models like CLIP (Radford et al., 2021) handle text and images by employing distinct tokenizers and encoders for each modality. However, those methods may encounter challenges in terms of efficiency and scalability when dealing with multimodal medical data. Additionally, maintaining separate tokenizers and encoders for each modality is inflexible. Inspired by Meta-transformer (Zhang et al., 2023a) and Imagebind (Girdhar et al., 2023), we employ a multimodal data tokenizer that converts various multimodal medical data into token embeddings. Due to extensive research on the conversion of 2D image data (e.g., X-ray and CT) and text data into tokens, there has been relatively less focus on tokenizing 3D tensor medical data (e.g., MRI and PET). Therefore, in the following section, we focus on transforming 3D tensor medical data into tokens.
@@ III-C1 Concept generation
A straightforward approach to generating a set of concepts for diagnosing neurological disorders is to leverage human expertise. Handcrafting a set of concepts offers better interpretability as they align with human perception and understanding (Wang et al., 2023). However, it is important to note that the process of annotating these concepts can be costly, as well as requiring rich medical expertise. To avoid those issues, we use GPT-4 (OpenAI, 2023) to automatically generate the diseases-related concepts. Specifically, by prompting GPT-4 with specific instructions, we can generate lists of concepts that are typically associated with diseases, e.g., Reduced Brain Metabolism: Individuals with AD typically exhibit slowed metabolic activity in the brain, particularly in the frontal and temporal lobes”. This approach reduces the annotation cost and leverages the capabilities of multimodal large models to generate concepts based on vast amounts of medical knowledge. More specifically, for $C$ categories in specific neurological disorder diagnoses, we generate $K$ relevant concepts for each category, which can be formalized as:
@@
where $\mathbf{GPT}$ denotes GPT-4 API (OpenAI, 2023), CLS denotes the name of corresponding category, and $\mathbf{X}_{\text{cpt}}$ denotes the generated text corresponding to $C\times K$ concepts.
@@ III-C2 Semantic similarity computation
where $\mathcal{C}_{i}$ is the set of concepts belongs to the category of this subject. Finally, we adjust the weight of the tokens and obtain the weight token embedding, i.e., $\tilde{t_{i}}=w_{i}t_{i}$. Note that, as the category of the sample is unknown, we select $\mathcal{C}_{i}$ belongs to which set of concepts with the highest weight during the inference.
@@ III-D Graph prompt learning
Neuroscience researchers (Rubinov and Sporns, 2010; Belmonte et al., 2004) have elucidated that the brain constitutes a complex graph structure, comprised of brain regions. The structural information of this graph, specifically the pattern of its connections, is crucial in the pathogenesis of neurological disorders. For example, the deterioration and abnormal connectivity are candidate biomarkers for Alzheimer’s disease (Pievani et al., 2014). Hence, the implementation of graph prompts is both reasonable and necessary for improving the ability of pre-trained multimodal large models to tackle the diagnosis of neurological disorders. To achieve this, we first construct the graph, and then extract embeddings from the constructed graph.
@@ III-D1 Graph construction
Considering that each token represent embedding of its corresponding patch (i.e., local brain region), we first regard the tokens as nodes in the graph. Then, we construct the edges/graph structural $\mathbf{A}$ based on semantic relationships between tokens. Notably, directly calculating the relationships between tokens based on token embeddings $\mathbf{T}$ may not be the optimal choice since token embeddings $\mathbf{T}$ may contain limited information about disease-related information. To solve this issue, we leverage the concept embeddings obtained in Section (III-C) as the bridge to learn the connections between tokens. With the the semantic similarity between tokens and concepts $\mathbf{S}$, the connection $\mathbf{a}_{i,j}$ between $i$-th token and $j$-th token is calculated by:
@@
where $\mathbf{S}_{i}$ is the semantic similarity between $i$-th token and all concepts, and $\tau$ denote temperature parameter. Intuitively, tokens belonging to similar concepts are more likely to be connected with a higher probability. This approach offers two advantages compared to directly calculating connection probabilities based on token embeddings. Firstly, it mitigates potential noise connection caused by irrelevant features in high-dimensional embedding $\mathbf{t}$. Secondly, the graph structure constructed based on concepts related to biomarkers and radiological characteristics of the disease is more meaningful in neuroscience. Based on this, we treat the constructed graph structure $\mathbf{A}$ as the prompt for pre-trained foundation modal.

### v2

@@ Abstract
Prompt learning has demonstrated impressive efficacy in the fine-tuning of multimodal large models to a wide range of downstream tasks. Nonetheless, applying existing prompt learning methods for the diagnosis of neurological disorder still suffers from two issues: (i) existing methods typically treat all patches equally, despite the fact that only a small number of patches in neuroimaging are relevant to the disease, and (ii) they ignore the structural information inherent in the brain connection network which is crucial for understanding and diagnosing neurological disorders. To tackle these issues, we introduce a novel prompt learning model by learning graph prompts during the fine-tuning process of multimodal models for diagnosing neurological disorders. Specifically, we first leverage GPT-4 to obtain relevant disease concepts and compute semantic similarity between these concepts and all patches. Secondly, we reduce the weight of irrelevant patches according to the semantic similarity between each patch and disease-related concepts. Moreover, we construct a graph among tokens based on these concepts and employ a graph convolutional network layer to extract the structural information of the graph, which is used to prompt the pre-trained multimodal models for diagnosing neurological disorders. Extensive experiments demonstrate that our method achieves superior performance for neurological disorder diagnosis compared with state-of-the-art methods and validated by clinicians.
@@ I Introduction
Figure 1: The flowchart of the proposed MMGPL consists of three modules i.e., multimodal data tokenizer (light blue block), concept learning (light green block), and graph prompt learning (light yellow block). First, MMGPL divides the multimodal medical data into multiple patches and project them into a shared embedding space (Sec. III-B). Second, MMGPL prompts the GPT-4 to generate disease-related concepts and further learn the weights of tokens based on the semantic similarity between tokens and concepts (Sec. III-C). Third, MMGPL learns a graph among tokens and extracts structural information to prompt the unified encoder (Sec. III-D). Finally, MMGPL obtains the output from the unified encoder and uses it to predict the label of the subject.
Firstly, only a small fraction of patches in neural data are related to the disease. Unlike natural image data, the pathological regions or the regions of interest (i.e., ROIs) in neural data typically only occupy a small portion of the entire data. For example, previous studies (Padurariu et al., 2012; Wang et al., 2006) have shown a significant correlation between the hippocampal area and Alzheimer’s disease. Recently, in the area of medical image segmentation, methods (Huang et al., 2023; Zhang et al., 2023c) utilize region-based prompts, such as bounding boxes and clicks, to guide the fundamental model focus on ROIs within the medical images. Despite the promising results obtained by region-based prompts, the aforementioned operations (e.g., bounding box and clicks) often rely on human interaction or object detection models which reduces the flexibility of these methods. Moreover, implementing the aforementioned operations becomes even more challenging when dealing with multimodality neural data, which may typically contains 3D tensor MRI data, time-series data, and functional connectivity (FC) networks data. Hence, it is challenging to discover useful patches when dealing with multimodality neural data.
@@ II-A Multimodal large models
Unlike traditional large language models such as GPT-3 (Brown et al., 2020) and BERT (Devlin et al., 2018), which can only deal with text data, multimodal large models (Radford et al., 2021; Tsimpoukelli et al., 2021; Liu et al., 2023; Driess et al., 2023) can integrate multiple modalities, such as language, vision, and audio, to perform a broad spectrum of tasks. For instance, CLIP (Radford et al., 2021) leverages a contrastive learning approach for joint pre-training on image-text pairs data. Besides, LLaVA (Liu et al., 2023) conducts image-text representation alignment followed by instruction tuning. Frozen (Tsimpoukelli et al., 2021) firstly proposed an approach that employs large language models in multimodal in-context learning. Similarly, Video-ChatGPT (Maaz et al., 2023) merges a video-adapted visual encoder with a large language model to handle video data. Recent works like Imagebind (Girdhar et al., 2023) and Meta-transformer (Zhang et al., 2023b) can learn a joint embedding across more than six different modalities.
In the medical domain, both large language models and multimodal large models have been recently explored for a wide range of medical tasks, including medical question answering and segmentation. For example, LLaVA-Med (Li et al., 2023) trains a vision-language conversational assistant research questions about medical images. MedSAM (Ma and Wang, 2023) fine-tune the segment anything model (SAM) (Kirillov et al., 2023) on medical image data for segmentation task. Moreover, Med-PaLM (Tu et al., 2023) and RedFM (Wu et al., 2023) interpret multimodal biomedical data and handle a diverse range of tasks, moving closer to generalist medical artificial intelligence models. To incorporate a large-scale multimodal dataset, Zheng et al. (2023) build up a large-scale diagnostic dataset comprising 39,026 cases and 192,675 medical images. They utilized this dataset to train a powerful multimodal medical foundation model. To ensure the trustworthiness of large models, Sun et al. (2024) introduce principles for different dimensions of trustworthiness. However, due to the complexity of multimodal neuroimaging data, there is still a lack of research and remain challenges in leveraging multimodal large models for diagnosing neurological disorders.
@@ II-B Prompt learning
Prompt learning has emerged as a new paradigm in fine-turning pre-trained large models for various downstream tasks. Compared to full fine-tuning, it achieves comparable performance without updating all parameters in large models. In the field of NLP, GPT-3 (Brown et al., 2020) has demonstrated strong generalization to downstream tasks by manually selecting prompt texts. To improve the construction of prompt texts, Prefix-tuning (Li and Liang, 2021) optimizes prompts through gradient-based fine-tuning. In computer vision, VPT (Jia et al., 2022) prompts the image for effectively using the pre-trained image encoder in the classification task. SAM (Kirillov et al., 2023) takes visual prompts by drawing points, boxes, and strokes on an image for image segmentation. Moreover, MaPLe (Khattak et al., 2023) utilizes multi-modal prompts to guide models in both image and text modalities. Besides, Miyai et al. (2024) apply prompt learning for few-shot out-of-distribution detection.
In the field of medicine, prompt learning aims to guide models by focusing their attention on relevant regions within complex medical datasets. For example, several researchers (Ma and Wang, 2023; Huang et al., 2023; Zhang et al., 2023c) have introduced prompts such as bounding boxes and clicks into the SAM model to adapt it for medical image segmentation tasks. However, these methods may not be suitable for diagnosing neurological disorders given that they rely on human interaction or object detection models, as well as they do not take into account the importance of brain connectivity in neurological disorders.
@@ II-C Graph neural network
In the past decade, Graph Neural Networks (GNNs) have been widely used in computer-aided diagnosis (Sun et al., 2020; Bessadok et al., 2022; Holzinger et al., 2021; Li et al., 2021). GNNs such as Graph Convolutional Network (GCN) (Kipf and Welling, 2017) and Graph Attention Network (GAT) (Veličković et al., 2017) leverage message passing mechanism to capture relationships between nodes and structure information in the graph. Based on this, GNNs have the capability to leverage the inherent relationships between brain regions to uncover patterns or biomarkers associated with neurological disorder. As a result, there has been a growing interest in applying GNNs in the field of neuroscience (Li et al., 2021). For example, Parisot et al. (2018) exploits the GCN for Alzheimer’s disease diagnosis in a population graph where the nodes in the graph denote the subjects and edges represent the similarity between subjects. Similarly, Kazi et al. (2019) proposes a GCN model with multiple filter kernel sizes on a population graph. Besides, Xu et al. (2023) leveraged graph structure information for biomedical synonym prediction.
In addition, brain graph is also frequently used in the diagnosis of neurological disorders, where the nodes correspond to anatomical brain regions and the edges depict functional or structural connections between these regions. For instance, Li et al. (2021) conduct an interpretable model on brain graph to determine the specific brain regions that are associated with a particular neurological disorder. Cai et al. (2022) propose a transformer-based geometric learning approach to handle multimodal brain graphs for brain age estimation. However, there is still a lack of research on how to effectively leverage graph in multimodal large models.
@@ III-A Preliminary and motivations
where $\text{LN}(\cdot)$ is layer normalization operation, $\mathbf{Z}^{(l-1)}$ is the output from the previous layer and the input token embeddings for $l=1$. In the MHSA layer, the attention mechanism computes multiple attention scores by performing dot products between the Query and Key vectors of the tokens, which represent dense interactions among tokens. A single-head attention operation can be expressed as follows:
@@
$\mathbf{P}_{\text{combined}}=\mathcal{F}_{\text{combine}}(\mathbf{P}_{\text{prompt}},\mathbf{P}),$ (4)
where $\mathcal{F}_{\text{combine}}$ denotes a fusion layer that can be a non-parametric operator ($\oplus$ or $+$) or parameterized layers. The combined token embeddings $\mathbf{P}_{\text{combined}}$ are subsequently processed by the transformer encoder, as delineated in Eq. 2. Based on this, previous methods proposed cross-modality prompt learning (Khattak et al., 2023) and deep prompt learning (Liu et al., 2022) to further improve its performance on specific tasks
@@ III-B Multimodal data tokenizer
Medical data is inherently multimodal, usually including modalities like MRI, PET, and FC in neurological disorder diagnosis. These diverse modalities provide complementary information, helping to gain a more comprehensive understanding and better analyze neurological disorders. However, these modalities often exhibit more complex data structures compared to natural images, such as 3D tensor medical data, brain connectivity graphs data, and time series data. Previous multimodal models like CLIP (Radford et al., 2021) handle text and images by employing distinct tokenizers and encoders for each modality. However, those methods may encounter challenges in terms of efficiency and scalability when dealing with multimodal medical data. Additionally, maintaining separate tokenizers and encoders for each modality is inflexible. Inspired by Meta-transformer (Zhang et al., 2023b) and Imagebind (Girdhar et al., 2023), we employ a multimodal data tokenizer that converts various multimodal medical data into token embeddings. Due to extensive research on the conversion of 2D image data (e.g., X-ray and CT) and text data into tokens, there has been relatively less focus on tokenizing 3D tensor medical data (e.g., MRI and PET). Therefore, in the following section, we focus on transforming 3D tensor medical data into tokens.
@@ III-C1 Concept generation
A straightforward approach to generating a set of concepts for diagnosing neurological disorders is to leverage human expertise. Handcrafting a set of concepts offers better interpretability as they align with human perception and understanding (Wang et al., 2023). However, it is important to note that the process of annotating these concepts can be costly, as well as requiring rich medical expertise. To avoid those issues, previous methods attempted to use GPT-4 (OpenAI, 2023) or other large language modal to generate concept texts. For example, (Yang et al., 2023) proposed to prompt the large language model (i.e., GPT-3) to generate candidate concepts. Based on this, (Yan et al., 2023) proposed to elicit medical knowledge from GPT-4 to build a set of concepts for interpretable medical image classifiers. Thus, building concepts from large language models (e.g., GPT-3 and GPT-4) has been success applied. In this study, we use GPT-4 to automatically generate the diseases-related concepts. Specifically, by prompting GPT-4 with specific instructions, we can generate lists of concepts that are typically associated with diseases, e.g., Reduced Brain Metabolism: Individuals with AD typically exhibit slowed metabolic activity in the brain, particularly in the frontal and temporal lobes”. This approach reduces the annotation cost and leverages the capabilities of multimodal large models to generate concepts based on vast amounts of medical knowledge. More specifically, for $C$ categories in specific neurological disorder diagnoses, we generate $K$ relevant concepts for each category, which can be formalized as:
@@
where $\mathbf{GPT}$ denotes GPT-4 API (OpenAI, 2023), CLS denotes the name of corresponding category, and $\mathbf{X}_{\text{cpt}}$ denotes the generated text corresponding to $C\times K$ concepts. In this study, the concept texts have been validated by clinicians to ensure they are correct.
@@ III-C2 Semantic similarity computation
where $\mathcal{C}_{i}$ is the set of concepts belonging to the category of this subject. Finally, we adjust the weight of the tokens and obtain the weight token embedding, i.e., $\tilde{t_{i}}=w_{i}t_{i}$. Note that, as the category of the sample is unknown, we select $\mathcal{C}_{i}$ belongs to which set of concepts with the highest weight during the inference.
@@ III-D Graph prompt learning
Neuroscience researchers (Rubinov and Sporns, 2010; Belmonte et al., 2004) have elucidated that the brain constitutes a complex graph structure, comprised of brain regions. The structural information of this graph, specifically the pattern of its connections, is crucial in the pathogenesis of neurological disorders. For example, deterioration and abnormal connectivity are candidate biomarkers for Alzheimer’s disease (Pievani et al., 2014). Hence, the implementation of graph prompts is both reasonable and necessary for improving the ability of pre-trained multimodal large models to tackle the diagnosis of neurological disorders. To achieve this, we first construct the graph, and then extract embeddings from the constructed graph.
@@ III-D1 Graph construction
Considering that each token represents embedding of its corresponding patch (i.e., local brain region), we first regard the tokens as nodes in the graph. Then, we construct the edges/graph structural $\mathbf{A}$ based on semantic relationships between tokens. Notably, directly calculating the relationships between tokens based on token embeddings $\mathbf{T}$ may not be the optimal choice since token embeddings $\mathbf{T}$ may contain limited information about disease-related information. To solve this issue, we leverage the concept embeddings obtained in Sec. III-C as the bridge to learn the connections between tokens. With the the semantic similarity between tokens and concepts $\mathbf{S}$, the connection $\mathbf{a}_{i,j}$ between $i$-th token and $j$-th token is calculated by:
@@
where $\mathbf{S}_{i}$ is the semantic similarity between $i$-th token and all concepts, and $\tau$ denote temperature parameter. Intuitively, tokens belonging to similar concepts are more likely to be connected with a higher probability. This approach offers two advantages compared to directly calculating connection probabilities based on token embeddings. Firstly, it mitigates potential noise connections caused by irrelevant features in high-dimensional embedding $\mathbf{t}$. Secondly, the graph structure constructed based on concepts related to biomarkers and radiological characteristics of the disease is more meaningful in neuroscience. Based on this, we treat the constructed graph structure $\mathbf{A}$ as the prompt for pre-trained foundation modal.